OpenStudy (anonymous):

Find the indicated outputs for f(x)=5x^2-2x f(0)= f(-1)= f(2)= My answer are: 1)1 2)25 3) 20

OpenStudy (saifoo.khan):

1 wrong. 2 wrong 3 wrong.

OpenStudy (anonymous):

@saifoo.khan you are killing me,lol

OpenStudy (saifoo.khan):

LOL.

OpenStudy (saifoo.khan):

f(x)=5x^2-2x f(0) = 5(0)^2 -2(0) = ?

OpenStudy (anonymous):

0

OpenStudy (saifoo.khan):

\[\checkmark\]

OpenStudy (saifoo.khan):

moving along..?

OpenStudy (anonymous):

I don't know,lol

OpenStudy (saifoo.khan):

lol, replace x with-1

OpenStudy (anonymous):

oh....f(-1)=5(-1)^2-2(-1)=

OpenStudy (saifoo.khan):

right.

OpenStudy (anonymous):

-20?

OpenStudy (saifoo.khan):

5+2 = 7. lol

OpenStudy (anonymous):

wait what....ummmm what?,lol

OpenStudy (saifoo.khan):

5(-1)^2-2(-1) 5(1) + 2 5 + 2 7

OpenStudy (anonymous):

But it's to the 2nd power?

OpenStudy (saifoo.khan):

-1 * -1 = +1

OpenStudy (anonymous):

That's true

OpenStudy (anonymous):

So, it's 7?

OpenStudy (saifoo.khan):

Yes momtoboys

OpenStudy (anonymous):

I was way off,lol

OpenStudy (saifoo.khan):

Haha, totally.

OpenStudy (anonymous):

hahaha....so (2) is 10?

OpenStudy (saifoo.khan):

Facepalm.

OpenStudy (anonymous):

haha Why??

OpenStudy (saifoo.khan):

how did u got 10? lol

OpenStudy (saifoo.khan):

teach me too.

OpenStudy (saifoo.khan):

haha

OpenStudy (anonymous):

I guessed...hehehe

OpenStudy (anonymous):

Ok let me think

OpenStudy (anonymous):

6?

OpenStudy (saifoo.khan):

No. :'(

OpenStudy (anonymous):

WHAT...don't cry, let me try again

OpenStudy (anonymous):

lol

OpenStudy (saifoo.khan):

lol

OpenStudy (anonymous):

8

OpenStudy (saifoo.khan):

is wall the right place to hit my head? lol

OpenStudy (anonymous):

Stop it I'm not that bad,lol

OpenStudy (saifoo.khan):

Haha, sorry i was just kidding. no hard feelings.

OpenStudy (saifoo.khan):

5(2)^2-2(2) 5(4) - 2(2) 20 - 4 16.

OpenStudy (anonymous):

That was my next answer

OpenStudy (saifoo.khan):

i have been tutoring for more than a year now.. i just keep math a little humours so people dont get bored and frustrated.

OpenStudy (saifoo.khan):

HAHA @next answer

OpenStudy (anonymous):

Well I like your humor thanks I need it...I hate math

OpenStudy (saifoo.khan):

Math is awesome.

OpenStudy (anonymous):

Nooooooo,lol

OpenStudy (saifoo.khan):

And thanks for not feeling bad about it. lol

OpenStudy (anonymous):

Not a problem!!

OpenStudy (saifoo.khan):

College Algebra is not that frustrating.

OpenStudy (anonymous):

Yes it is

OpenStudy (saifoo.khan):

NOOOOOOOOOOOOOOO!